OpenStudy (anonymous):

How should I begin to calculate the following limit? \[\lim_{x \rightarrow +\infty} (\frac{x+4}{x+2})^x\] Does it have relation with the orders of both numerator and denominator? So the answer to the question should be 1? Or is it wrong?

OpenStudy (anonymous):

\[\lim_{x \rightarrow +\infty} (\frac{x+4}{x+2})^x\]

OpenStudy (anonymous):

take the log. then compute the limit, then exponentiate. by which i mean if your answer to the limit of the log is 3, your actual answer will be \[e^3\]

OpenStudy (anonymous):

so step one is \[x\ln(\frac{x+4}{x+2})\] which looks like infinity times 0

OpenStudy (anonymous):

step two rewrite as \[\frac{\ln(\frac{x+4}{x+2})}{\frac{1}{x}}\]

OpenStudy (anonymous):

standard trick when you are faced with \[0\times \infty\]

OpenStudy (anonymous):

now use l'hopital to find your limit

OpenStudy (zarkon):

the final answer is \[e^2\]

OpenStudy (anonymous):

not so easy now is it? but that is ok just need the derivative of top and bottom derivative of \[\frac{1}{x}\] is \[-\frac{1}{x^2}\]

OpenStudy (anonymous):

derivative of \[\ln(\frac{x+4}{x+2})\] is \[-\frac{2}{x^2+6x+8}\]

OpenStudy (anonymous):

so bring up the denominator and you get \[\frac{2x^2}{x^2+6x+8}\] take the limit as x goes to infinity with your eyeballs and get limit is 2 so answer is... what zarkon said!

OpenStudy (anonymous):

all the gory details are there so i hope it is clear, more or less

OpenStudy (anonymous):

Thank you very much, satellite73! Very clear explanation!

OpenStudy (zarkon):

if you know that \[\lim_{x\to\infty}\left(1+\frac{r}{x}\right)^{tx}=e^{rt}\] then you can do the following \[\lim_{x \rightarrow +\infty}\left (\frac{x+4}{x+2}\right)^{x}\] \[=\lim_{x \rightarrow +\infty}\left (\frac{x+2+2}{x+2}\right)^{x}\] \[=\lim_{x \rightarrow +\infty}\left (\frac{x+2}{x+2}+\frac{2}{x+2}\right)^{x}\] \[=\lim_{x \rightarrow +\infty}\left (1+\frac{2}{x+2}\right)^{x}\] \[=\lim_{x \rightarrow +\infty}\left (1+\frac{2}{x+2}\right)^{x+2-2}\] let \[w=x+2 \text{ then as }x\to\infty, w\to\infty\] \[=\lim_{w \rightarrow +\infty}\left (1+\frac{2}{w}\right)^{w-2}\] \[=\lim_{w \rightarrow +\infty}\left (1+\frac{2}{w}\right)^{w}\left (1+\frac{2}{w}\right)^{-2}\] \[=e^2\cdot 1=e^2\]

OpenStudy (anonymous):

wow. i guess you can do that. it is good to know this way too.

OpenStudy (zarkon):

satellite73 technique is needed to verify that \[\lim_{x\to\infty}\left(1+\frac{r}{x}\right)^{tx}=e^{rt}\]

OpenStudy (zarkon):

but once you have it ...you might as well use it ;)

OpenStudy (anonymous):

Yeah, thanks Zarkon for this another way of solving it!

OpenStudy (anonymous):

@zarkon no, not really. in fact that \[\lim_{n\rightarrow \infty}(1+\frac{x}{n})^{n}\] is as good a definition of \[e^x\] as any

OpenStudy (zarkon):

if that is your definition of e^x then sure otherwise it needs some kind of proof

myininaya (myininaya):

gj you guys i was kindof thinking of zarkon's way before i scrolled down and see his answer i like his answer best :)

OpenStudy (anonymous):

no you don't! you are just trying to be mean to me. i bet you $1,000,000,000 that you would have done just what i did. take the log etc. so don't try to play me young lady. i have a whole stalk of sensei

myininaya (myininaya):

lol